OpenStudy (anonymous):

The numbers 20,21,29 form a Pythagorean triple. Which other set of numbers is a Pythagorean triple?

OpenStudy (anonymous):

12-5-13

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

20, 21, and 29 are a Pythagorean triple, because when you square 20 and 21, the sum of the two is equal to 29 squared. You must give another example of this.

OpenStudy (anonymous):

12,5,13

OpenStudy (anonymous):

square 12 and 5 and it will be equal to 13 square

OpenStudy (anonymous):

\[20^{2} + 21^{2} = 29^{2}\] is the same as 400 + 441 = 841

OpenStudy (anonymous):

basically a Pythagorean triplet is a combination of whole numbers with no fractional part

OpenStudy (anonymous):

hopw you get it now

OpenStudy (anonymous):

the numbers form a Pythagorean triple. Which other set of numbers is a Pythagorean triple? choices are A.21,22,30 B.40,42,58 C.60,66,87 D.70,84,116

OpenStudy (anonymous):

So using Hashir's example: \[12^{2} + 5^{2} = 13^{2}\] is the same as 144 + 25 = 169 which is true. So 12, 5, 13 is a Pythagorean Triplet

OpenStudy (anonymous):

b

OpenStudy (anonymous):

the numbers 20,2129 form a Pythagorean triple. Which other set of numbers is a Pythagorean triple? choices are A.21,22,30 B.40,42,58 C.60,66,87 D.70,84,116

OpenStudy (anonymous):

its an easy one notice that B is the multiple of the given Pythagorean triple ....... so it satisfies the situation

OpenStudy (anonymous):

is is say that 12-5-13 is a pythagorean triplet than i can also say that 24-10-26 is a pythagorean triplet

OpenStudy (anonymous):

is is= i can (TYPO)

OpenStudy (anonymous):

B is the correct answer. 40^2 + 42^2 = 58^2 is the same as 1600 + 1764 = 3364 which is true.